### v1

# Initialization Is Critical to Whether Transformer fits Composite Function by Inference or Memorizing

Abstract
Transformers have shown impressive capabilities across various tasks, but their performance on compositional problems remains a topic of debate. In this work, we investigate the mechanisms of how transformers behave on unseen compositional tasks using anchor functions. We discover that the parameter initialization scale plays a critical role in determining whether the model learns inferential solutions, which capture the underlying compositional primitives, or symmetric solutions, which simply memorize mappings without understanding the compositional structure. By analyzing the information flow and vector representations within the model, we reveal the distinct mechanisms underlying these solution types. We further find that inferential solutions exhibit low complexity bias, which we hypothesize is a key factor enabling them to learn individual mappings for single anchors. Building upon our understanding of these mechanisms, we can predict the learning behavior of models with different initialization scales when faced with data of varying inferential complexity. Our findings provide valuable insights into the role of initialization scale in shaping the type of solution learned by transformers and their ability to learn and generalize compositional functions.
1 Introduction
Large-scale transformers demonstrate unprecedented capabilities (Achiam et al., 2023; Teubner et al., 2023; Choi et al., 2021; Brown et al., 2020), even noted as "sparks of AGI" (Bubeck et al., 2023). However, their performance on compositional tasks, which are considered a key property of human cognition (Marcus, 2003; Smolensky et al., 2022), has long been the subject of intense debate, especially in terms of out-of-distribution (OOD) generalization ability. Furthermore, the mechanisms underlying how large language models accomplish compositional tasks remain unclear. This raises critical open questions about how to faithfully interpret transformers’ capabilities on compositional tasks. Can transformers learn the underlying compositional primitives within the data, or do they merely learn input-output mappings? When transformers produce incorrect outputs on compositional tasks, do their responses follow any discernible patterns or logic?
In this work, we use anchor functions (Zhang et al., 2024) to investigate how transformers generalize to unseen compositional tasks. In our setup, each sequence contains a composite anchor pair, a key item, and noise items unrelated to the output (right side of Fig. 1(a)). The single anchors are specific tokens (i.e., 1, 2, 3, 4), each corresponding to a particular arithmetic operation (left side of Fig. 1(a)). A composite function is defined as the sequential application of two single anchor functions, i.e., the corresponding arithmetic operation, to a key item. For example, given a key item $x$ and a composite anchor pair (1, 2), the composite function would output $(x+5)+1$. As shown in the middle part of Fig. 1(a), during training, 14 out of the 16 possible composite anchor pairs are assigned inferential mappings, meaning the target output is consistent with the result of applying the composite function. One pair (3, 4) is assigned a non-inferential mapping, where the target output deviates from the result of the composite function. The remaining pair (4, 3) is held out as an unseen task. After training, there are three possible solutions the model could learn for the unseen pair (4, 3): an inferential solution, a symmetric solution matching the non-inferential seen pair (3, 4), or other non-inferential solutions (Fig. 1(b)). The central question we aim to answer is: which of these three types of solutions will the model learn for the unseen anchor pair (4, 3)?
Figure 1: Experimental setup and possible solutions and mechanisms for the unseen composite anchor pair (4, 3). (a) Data generation: Left: The single anchors (1, 2, 3, 4) correspond to specific arithmetic operations. Middle: During training, 14 out of the 16 possible composite anchor pairs are assigned inferential mappings, one pair (3, 4) is assigned a non-inferential mapping, and the remaining pair (4, 3) is held out as an unseen task (does not appear in the training set). Right: The input sequences comprise a composite anchor pair, a key item preceding the anchor pair, and other unrelated items. The question mark indicates the output for the unseen anchor pair (4, 3), which depends on the learned solution. (b) Possible solutions and mechanisms for the unseen anchor pair (4, 3): After training, the model may learn an inferential solution, a symmetric solution matching the non-inferential seen pair (3, 4), or other non-inferential solutions not shown in the figure. Two potential mechanisms for task generalization are illustrated: learning the symmetric structure (Mechanism 1) or composing the inferred single anchor mappings (Mechanism 2).
Based on the above experimental setup, we find that the model exhibits two distinct phases of solutions depending on the parameter initialization scale with different model depths. With small initialization scales, the model tends to learn inferential solutions on the unseen composite anchor pair (4, 3). In contrast, with larger initialization scales, the model is more likely to learn symmetric solutions on anchor pair (4, 3) that simply match the output of its symmetric anchor pair (3, 4). These findings suggest that the initialization scale plays a crucial role in determining the type of solution learned by the model.
To gain insights into the mechanisms underlying these different solution types, we analyze the information flow and vector representations within the model. We find that symmetric solutions directly combine anchor information without capturing the underlying compositional primitives, while inferential solutions learn individual mappings for each single anchor and compose them to generate the final output. We hypothesize that the model’s ability to learn compositional primitives is influenced by its complexity, determined by the initialization scale. Inferential solutions exhibit low complexity, with weights condensing in a few directions and input tokens arranged according to their numerical value in the embedding space, while symmetric solutions show no obvious condensation and lack clear structural features, indicating higher complexity.
Building upon our understanding of these mechanisms, we predict and experimentally validate that models with small initialization scales require more training steps to fit data with larger inferential complexity, defined as the number of composite anchor pair groups that do not satisfy the inferential mapping, while models with large initialization scales use similar training steps regardless of inferential complexity.
In summary, our work highlights the crucial role of initialization scale in shaping the solutions learned by transformers on compositional tasks. These insights can inform the design of future transformer models, enabling them to better capture compositional structures and generalize to unseen tasks, ultimately leading to improved performance on compositional problems.
2 Related Work
Challenges of transformers in compositional tasks.
Recent advancements in LLMs have showcased remarkable capabilities across a wide range of tasks, often surpassing human performance (Wei et al., 2022; Fu et al., 2022). However, despite their impressive performance on single-step reasoning tasks (Srivastava et al., 2022), transformers face significant challenges when it comes to effectively combining multiple steps to solve compositionally complex problems (Hupkes et al., 2018; Csordás et al., 2021, 2022; Yun et al., 2022; Lepori et al., 2023; Okawa et al., 2023; Dziri et al., 2024). Ramesh et al. (2023) show that transformers trained to directly compose capabilities struggle to generalize to out-of-distribution (OOD) tasks. Liu et al. (2022) suggest that shallow transformers learn shortcuts during training, leading to poor OOD generalization. Numerous studies have explored various approaches to address these limitations, such as encouraging explicit reasoning step generation within a single generation (Wei et al., 2022), leveraging LLMs to generate reasoning steps iteratively (Creswell et al., 2022; Creswell and Shanahan, 2022). Despite these efforts, achieving complete mastery in compositional tasks remains a significant challenge for vanilla transformers. Zhang et al. (2024) propose a type of benchmark functions called anchor functions for studying language models. Building upon the compositional task settings in anchor functions, our work investigates the solutions and underlying mechanisms of models under different initialization scales.
Parameter initialization of neural networks.
The parameter initialization of the network is important to determine the final fitting result of the network (Jacot et al., 2018; Arora et al., 2019; Zhang et al., 2019; Mei et al., 2018; Rotskoff and Vanden-Eijnden, 2018; Chizat and Bach, 2018; Sirignano and Spiliopoulos, 2020; Williams et al., 2019; E et al., 2020). Luo et al. (Luo et al., 2021), Zhou et al. (Zhou et al., 2022) mainly identify the linear regime and the condensed regime for two-layer and three-layer wide ReLU NNs, respectively. A series of works suggest that the condensed regime improves the generalization ability of models (Zhang et al., 2023; Zhang and Xu, 2024; Zhang et al., 2021, 2022). Recent studies have also investigated the impact of initialization on the training of large language models (LLMs) (Liu et al., 2020; Zhu et al., 2021; Trockman and Kolter, 2023; Zhang et al., 2019; Huang et al., 2020; Wang et al., 2024). These works primarily focus on how the initialization scale affects training stability and other aspects of LLM performance. In our work, we find that different initialization scales can influence a model’s ability to memorize and infer data on the compositional task.
3 Definitions
3.1 Two-anchor composite function
A two-anchor composite function $f(X):\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{d}$ is defined as
$f(\bm{x}_{1},\ldots,\bm{x}_{n})=g\left(g(\bm{x}_{i-1};\bm{x}_{i});\bm{x}_{i+1}\right),\quad\text{where}\quad\bm{x}_{i},\bm{x}_{i+1}\in A.$ (1)
Here, the input sequence $X=(\bm{x}_{1},\ldots,\bm{x}_{n})$ comprises $n$ tokens, each of dimension $d$. An anchor set $A=\{\bm{a}_{1},\bm{a}_{2},\ldots,\bm{a}_{J}\}$ is designated, with each token $\bm{a}_{k}\in A$ corresponds to a function $g(\bm{x};\bm{a}_{k})$. In each $X$, one and only one pair of two consecutive elements belong to $A$, such as $\bm{x}_{i},\bm{x}_{i+1}\in A$. To simplify notation, we denote the two-anchor composite function as $f(\bm{x}_{i-1};\bm{x}_{i},\bm{x}_{i+1})$ to emphasize the composite anchor pair $(\bm{x}_{i},\bm{x}_{i+1})$ and key item $\bm{x}_{i-1}$.
For example, consider a simple case where $d=1$ and the anchor set $A=\{1,2,3,4\}$. Each anchor token corresponds to a specific function:
$g(x;1)=x+5,\quad g(x;2)=x+1,\quad g(x;3)=x-2,\quad g(x;4)=x-8.$ (2)
Given $X=(23,1,2,43,46,74,44,72)$, the two-anchor composite function would output:
$f(X)=f(23;1,2)=g(g(23;1);2)=(23+5)+1=29.$ (3)
3.2 Data Generation
In this work, we construct the input dataset using four anchors (i.e., 1, 2, 3, 4) and 80 key items (i.e., 20-99). The four anchors comprise 16 composite anchor pairs. We select a subset or all of the composite anchor pairs to construct the training dataset based on the task requirements.
By default, i.e., with the inferential mapping defined below, the target is the output of the input data processed by the two-anchor composite function. Our definition of single-anchor function is consistent with the definition in Equ. (2).
3.3 Mapping Type of a Composite Anchor Pair
For a composite anchor pair ($a_{1}$, $a_{2}$), the type of its mapping $\mathcal{M}_{(a_{1},a_{2})}(\cdot)$ can be classified as follows.
Inferential mapping. The designated target mapping of a composite anchor pair ($a_{1}$, $a_{2}$) is consistent with the two-anchor composite function, i.e., $\mathcal{M}_{(a_{1},a_{2})}(x)=f(x;a_{1},a_{2})$.
Non-inferential mapping. The designated target mapping of a composite anchor pair ($a_{1}$, $a_{2}$) is inconsistent with the two-anchor composite function, i.e., $\mathcal{M}_{(a_{1},a_{2})}(x)\neq f(x;a_{1},a_{2})$.
Symmetric mapping. The designated target mapping of a composite anchor pair ($a_{1}$, $a_{2}$) is consistent with the mapping of its symmetric composite anchor pair ($a_{2}$, $a_{1}$), i.e., $\mathcal{M}_{(a_{1},a_{2})}(x)=\mathcal{M}_{(a_{2},a_{1})}(x)$.
We refer to a model as an inferential (non-inferential, symmetric) solution if it tends to output the mapping corresponding to the inferential (non-inferential, symmetric) mapping for the studied composite anchors.
3.4 Generalization
The division of the dataset naturally leads to the following two concepts of generalization:
Generalization on data. Generalization on data primarily relies on the test set, which is divided according to key items. In this test data, all anchor pairs (i.e., the task) are seen in the training set.
Generalization on task. Generalization on task primarily depends on the unseen anchors, i.e., anchors that do not appear in the training set, with a designated target mapping.
3.5 Model Architecture and Basic Experimental Setups
The input sequence is represented as a one-hot vector $X^{\mathrm{in}}$. The word embedding $X^{\mathrm{em}}$ and the input to the first transformer block $X^{(1)}$ is calculated as:
$X^{\mathrm{em}}=X^{\mathrm{in}}W^{\mathrm{em}},X^{(1)}=X^{\mathrm{em}}+X^{\mathrm{pos}},$ (4)
where $X^{\mathrm{pos}}$ is the positional vector. For the $l$-th layer, the $Q,K,V$ are defined as:
$Q^{(l)}=X^{(l)}W^{Q(l)},\quad K^{(l)}=X^{(l)}W^{K(l)},\quad V^{(l)}=X^{(l)}W^{V(l)}.$ (5)
The attention matrix $\mathrm{Attn}^{(l)}$ and its subsequent output $X^{\mathrm{qkv}(l)}$ for the $l$-th layer is computed as:
$\mathrm{Attn}^{(l)}=\mathrm{softmax}\left(\frac{Q^{(l)}K^{(l)T}}{\sqrt{d_{k}}}\right)\text{ (with mask)},\quad X^{\mathrm{qkv}(l)}=\mathrm{Attn}^{(l)}V^{(l)}.$ (6)
The output of the $l$-th attention layer is obtained as:
$X^{\mathrm{ao}(l)}=\text{LN}(X^{(l)}+X^{\mathrm{qkv}(l)}W^{\mathrm{attn},l}),\quad X^{(l+1)}:=X^{\mathrm{do}(l)}=\text{LN}(\text{FNN}(X^{\mathrm{ao}(l)})+X^{\mathrm{ao}(l)}).$ (7)
The final output is obtained by projecting the output of the last layer $X^{\mathrm{do}(L)}$ using a linear projection layer, followed by a softmax operation and argmax to obtain the predicted token.
It is worth noting that the model architecture described above is for a single-head attention model.
For the basic experimental setups, we use cross-entropy loss on the last token of the sequence and optimize the model using Adam with weight decay. The specific hyperparameters and training details are provided in Appendix A.
4 Two Phases of Solutions for Composite Functions
In this section, we investigate the mechanisms of how a transformer network model fits the training data and its generalization.
Experimental Setup. The anchor pair (4,3) is held out as an unseen pair during the training, while pair (3,4) is assigned as non-inferential mapping, i.e., we set the designated target mapping of ($\cdots$, $x$, 3, 4, $\cdots$) to $x-6$ (the inferential mapping is $x-10$), seen in the training set. The other 14 anchor pairs are set as inferential mappings seen in the training set. Three possible solutions for pair (4,3) may be learned after enough training: i) inferential solution based on the 14 anchor pairs with inferential mapping, ii) symmetric solution based on the anchor pair (3,4), iii) other non-inferential solutions. See illustration in Fig. 1(b).
Figure 2: Phase diagram of generalization performance on unseen anchor (4, 3). (a) The model’s test accuracy based on the symmetric mapping. (b) The model’s test accuracy based on the inferential mapping. The abscissa represents the initialization rate $\gamma$, which corresponds to the standard deviation $(1/d_{\mathrm{in}})^{\gamma}$ of a normal distribution with a mean of 0 used for parameter initialization. The ordinate represents the depth of the transformer model. The shadow zones indicate the test accuracy on seen anchors is less than 90%.
An important and interesting question is: which solution is learned for unseen pair (4,3)? Through experiments with different parameter initialization scales and model depths, we find that the model use different mechanisms to fit training data, leading to different solutions for unseen pair (4,3). To see if the model learns the symmetric solution for the unseen anchor pair (4,3) during the training, we can plot the phase diagram of the model accuracy for the anchor pair (4,3) by assuming the ground truth as the symmetric mapping, against the parameter initialization scale and model depth. As shown in Fig. 2(a), the yellow zones with the large initialization scale (small $\gamma$) indicate $100\%$ test accuracy, i.e., the model learns the symmetric solution for anchor pair (4,3). Similarly, Fig. 2(b) assumes the ground truth as the inferential mapping. For small initialization scale (large $\gamma$), the model tends to learn the inferential solution for the anchor pair (4,3). When the initialization scale is too large and the model is deep (shadow zone), the model learns a non-inferential solution, which generalizes badly even on seen anchors.
We can conclude that the model tends to learn the inferential solution with small initialization while the symmetric solution with large initialization (but not very large). Note that this result is rather stable as we obtain these two phase diagrams over many trials as follows. For a specific random seed, models with different initializations and depths are trained using 9 learning rates (uniformly sampled in the interval [$1\times 10^{-4}$, $3\times 10^{-4}$]), and the highest accuracy for the inferential and symmetric mappings is selected. We use 3 random seeds for the experiments and take the average of the highest accuracy obtained from the three seeds.
5 Mechanisms of Models in Two Phases
In this section, we analyze mechanisms underlying different learning solutions including: i) information flow via attention matrix; ii) information representation in vector space; iii) model complexity form different perspectives. A key message is as follows. A model with small initialization needs to gradually increase its capacity by enlarging the parameter scale, therefore, to use the least optimization cost to fit the training data, the model only needs to learn four elementary anchor functions to obtain inferential mappings. As the initialization scale gets larger, it tends to learn ten symmetric mappings, and memorize the mapping between anchor-key item pairs and their outputs.
5.1 Information Transmission and Fusion Mechanisms
In this section, we study the information transmission and fusion mechanisms occurring in the attention matrix through information flow analysis. As shown in Fig. 3 (a,c), we present the information flow of two-layer networks corresponding to two solutions. We use the same input sequence $X=(35,63,99,4,3,38,42,36,62)$ to test the output of the key item $x_{3}=99$ under the influence of the composite anchor pair $(x_{4},x_{5})=(4,3)$. The thickness of the line connecting the $j$-th token in Layer $l$ and the $k$-th token in Layer $l+1$ represents the value of the attention matrix $\mathrm{Attn}^{(l)}$ at position ($k$, $j$). For ease of observation, we manually annotate the information flow that significantly contributes to the output of the last token using different colors.
Figure 3: (a, c) Information flow in the two-layer networks of symmetric and inferential solutions. The input sequence shown in the figure represents the test sample, with key items and anchor positions annotated. For each layer’s attention matrix, we illustrate the mechanisms of information transmission and fusion through the information flow. The thickness of the line represents the value of the attention matrix $\mathrm{Attn}^{(l)}$. We use different colors to mark the key item and the two single anchors, and highlight the attention connections that significantly contribute to the final output. The final output sequence represents the model’s output. (a) Symmetric solution. (c) Inferential solution. (b) T-SNE visualization of vectors $X^{\mathrm{ao}(1)}$ of 10,000 input sequences with different anchor-key item pairs.
For symmetric solutions, we find that the model’s attention mechanism plays a key role in enabling the learning of consistent mappings for symmetric anchor pairs. In the first layer of the network, the attention mechanism combines the information from the two anchors and places it in the last token of the sequence. This allows the model to learn identical embeddings for symmetric anchor-key pairs, such as (3, 4) and (4, 3), as illustrated in Fig. 3(a). In the second layer, the information of the key item is broadcast to the last token and combined with the anchor information obtained from the first layer by the residual connection. This enables the model to produce the final output based on the combined information from the anchors and the key item.
To verify the symmetric nature of the learned representations after the first-layer attention, we visualize the vectors $X^{\mathrm{ao}(1)}$ of 10,000 input sequences with different anchor-key item pairs using t-SNE. Fig. 3(b) shows that symmetric anchor-key pairs (e.g., (3, 4) and (4, 3)) are clustered together in the $X^{\mathrm{ao}(1)}$ vector space, confirming that the model learns to map them to similar representations.
For inferential solutions, the information transmission mechanism is different. As shown in Fig. 3(c), in the first layer, the key item information is moved to the positions of the two anchors, and each anchor is combined with the key item information separately. This allows the model to learn individual mappings for each single anchor. In the second layer, the anchor tokens (now containing the combined information from the anchors and the key item) are broadcast to the last token and combined to produce the final output. This composition of single anchor mappings enables the model to learn inferential solutions that generalize to unseen composite anchors.
5.2 Divergence in Fused Vector Representations across Two Phases
To further investigate the mechanistic differences between symmetric and inferential solutions, we examine the divergence in vector representations after the fusion of anchor and key item information. Specifically, for different anchor-key item combinations, we explore the cosine similarity between the output vectors $X^{\mathrm{ao}(2)}_{[9,:]}$, i.e., the last token of the output vectors of the second attention layer, under the two mechanisms. Let $\bm{v}_{(x;a_{1},a_{2})}$ denote the output vector $X^{\mathrm{ao}(2)}_{[9,:]}$ for a sequence with key item $x$ and anchors $(a_{1},a_{2})$ [^1] For simplicity, we assume that the position of the anchor-key item combination and other items do not significantly affect the output vector..
As shown in Fig. 4 (a, b), the heatmaps display the cosine similarity between output vectors for different key item and anchor combinations. Both the $x$-axis and $y$-axis represent the values of key items, ranging from 30 to 40. The lower triangle of the heatmap corresponds to the cosine similarity matrix between $\bm{v}_{(x;3,3)}$ and $\bm{v}_{(y;2,2)}$, where $x$ and $y$ are the key items at the corresponding positions on the $x$-axis and $y$-axis, respectively. Similarly, the upper triangle matrix represents the similarity matrix between $\bm{v}_{(x;1,2)}$ and $\bm{v}_{(y;1,3)}$. The red boxes highlight the positions where the inferential targets are the same, for example, the red highlight in the first column in Fig. 4(a) indicates $f(36;3,3)=f(30;2,2)$.
For inferential solutions, and for any key items $x_{1}$, $x_{2}$, anchors $(a_{1,1},a_{1,2})$, $(a_{2,1},a_{2,2})$, if $f(x_{1};a_{1,1},a_{1,2})=f(x_{2};a_{2,1},a_{2,2})$, then the vectors $\bm{v}_{(x_{1};a_{1,1},a_{1,2})}$ and $\bm{v}_{(x_{2};a_{2,1},a_{2,2})}$ are nearly parallel. In contrast, for symmetric solutions, the pairwise similarity between vectors $\bm{v}_{(x;a_{1},a_{2})}$ is always low even for those with the same output. This indicates that the last FNN layer of the symmetric solution memorize all possible projections from $X^{\mathrm{ao}(2)}_{[9,:]}$ to the output of the decoder layer $X^{\mathrm{do}(2)}_{[9,:]}$, while the last FNN for the inferential solution only needs to learn much fewer projections.
Figure 4: Cosine similarity heatmaps. Each axis represents a selected anchor, with the value on the coordinate axis representing the value of the key item. Color indicates the cosine similarity between specific vectors, which are defined within the respective subplot. The red boxes highlight the positions where the target outputs obtained by the anchors on the $x$-axis and $y$-axis are the same for the corresponding key items. (a, b) Color indicates the cosine similarity between the last token of $X^{\mathrm{ao}(2)}$ of different anchor-key item pairs. (a) Inferential solution. (b) Symmetric solution. (c) Color indicates the cosine similarity between $V^{(2)}$ rows corresponding to the first anchor’s position across different anchor-key pairs for inferential solutions.
The core of learning inferential solutions lies in the model’s ability to infer the operation of each single anchor. To further verify the alignment between the single anchor operations in the inferential mechanism and the defined single-anchor function $g$, we investigate the vector representations after the fusion of the first anchor and the key item. From the information flow diagram (Fig. 3(c)), the information from the two anchors and the key item is fused through the attention mechanism in the second layer. Therefore, we focus on the vector representation of the second layer Value matrix $V^{(2)}$ at the position of the first single anchor. Similar to the approach in Fig. 4(a, b), we examine the structural characteristics of the output $\bm{u}_{(x;a_{1})}$ at the position of the first anchor of $V^{(2)}$ for the sequence with key item $x$ and the first anchor $a_{1}$. The red boxes highlight the positions where the targets obtained by the single anchors $a_{1}$ and $a_{2}$ on the $x$-axis and $y$-axis are the same for the corresponding key items $x_{1}$ and $x_{2}$, i.e., $g(x_{1};a_{1})=g(x_{2};a_{2})$. As shown in Fig. 4(c), for single anchor mappings, the information fusion process of single anchors and key items during the inferential solution’s information fusion and the explicitly defined single anchor operations align well. This provides stronger evidence for the model’s ability to learn inferential solutions.
5.3 Model Complexity: A Key Factor in Phase Transitions
Large initialization scale endows the model with large complexity with training data with changing parameters slightly, such as the case of fully-connected neural networks (FNN) in linear regime (e.g., Neural tangent kernel) (Luo et al., 2021; Jacot et al., 2018). However, the model with small initialization with inadequate complexity often gradually increases its complexity during the training. For example, neurons in the same layer of FNNs with small initialization would first condense in very few directions (Zhou et al., 2021), followed by condensing in more and more directions (Luo et al., 2021; Zhang et al., 2021, 2022), which can regularize FNNs to fit data with as few condensation directions as possible. Therefore, we investigate the degree of parameter condensation in models with inferential or symmetric solutions.
Figure 5: (a, b) Cosine similarity of neurons in the $W^{Q(1)}$ matrix. (a) Inferential solution with small initialization. (b) Symmetric solution with large initialization. (c, d) Visualization of the embedding space using PCA for different initialization scales. (c) Inferential solution with small initialization. The embedded tokens seem to form arithmetic sequences with common differences of 3 (red arrow) and 4 (blue arrow) along the two directions. (d) Symmetric solution with large initialization.
Condensation. We examine the condensation of matrix $W^{Q(1)}$. The similarity between the $i$-th and the $j$-th neurons input weight is calculated by cosine similarity, i.e., $\frac{W^{Q(1)}{[i,:]}\cdot W^{Q(1)}{[j,:]}}{||W^{Q(1)}{[i,:]}||_{2}||W^{Q(1)}{[j,:]}||_{2}}$. As shown in Fig. 5, for the inferential solution (Fig. 5(a)), the neuron weights condense in a few directions, suggesting low complexity within the model. In contrast, for the symmetric solution (Fig. 5(b)), there is no obvious condensation of neuron parameters, indicating high complexity within the model.
However, even with small initialization, not all parameters exhibit significant condensation. For the word embedding matrix $W^{\mathrm{em}}$, to distinguish the meanings of different tokens, different neuron weights have different directions, i.e., neurons do not condense. Nevertheless, the parameter matrix still exhibits a clear tendency towards low complexity.
Structural Features of the Word Embedding Matrix. To gain further insights into the learning mechanisms of inferential solutions, we analyze the structure of the model’s embedding matrix. We visualize the embedding space using Principal Component Analysis (PCA). As shown in Fig. 5(c,d), for small initialization scales (Fig. 5(c)), we observe a clear ordinal structure in the embedding space, with the embeddings of the input tokens arranged according to their numerical value. This also suggests a low-complexity tendency in the word embedding matrix, requiring the model to capture the relative ordinal relationships between different tokens. However, this ordinal structure is not present for large initialization scales (Fig. 5(d)).
It is worth noting that the relative ordinal relationship in the inferential solution is not a simple numerical magnitude relationship of the corresponding tokens. We observe that this ordinal relationship may originate from the definition of the four single anchors, where the differences between the operations of any two single anchors can be obtained by the addition of the basic elements 3 and 4. This arrangement leads to the numbers being ordered with intervals of 3 (red arrow) and 4 (blue arrow) from two directions in the embedding space. To further verify the low complexity bias of word embedding matrices, we visualize the eigenvalues of the covariance matrix of the embedding vectors, the results are shown in Appendix B.
6 Prediction of mechanisms underlying different solutions
As the data complexity increases, neural networks usually need more training steps to fit the data (Arpit et al., 2017; Xu et al., 2019, 2020; Rahaman et al., 2019). We define the number of composite anchor pair groups [^2] Two anchor pairs with symmetrical relationships are considered to be in the same group. that do not satisfy the inferential mapping as the inferential complexity.
It is anticipated that the model of small initialization, i.e., fitting data with as low complexity as possible, should use more training steps to fit data with larger inferential complexity, while the model with large initialization, i.e., easily memorizing ten mappings of the symmetric solution, should use similar training steps to fit data with different inferential complexity.
Figure 6: Training speed with different initialization scales when facing different data complexity. Abscissa: the model depth. Ordinate: the number of epochs required for the model’s training loss to reach $5\times 10^{-2}$. Color: the number of noise groups. We conduct 9 random trials for each setting of different initializations, depths, and data inferential complexities, and take the average of the required number of epochs. (a) Small initialization. (b) Large initialization.
To measure the training steps required to fit data, we use the number of epochs required for the model’s training loss to reach $5\times 10^{-2}$. We conduct 9 random trials for each setting. As shown in Fig. 6, models that tend to learn data with inferential solutions indeed use more training steps for data with larger inferential complexity (Fig. 6(a)), while models that tend to learn data with symmetric solutions use roughly the same training steps for data with different inferential complexity (Fig. 6(b)). This prediction convincingly backs up our analysis of why different initialization scales can lead to different solutions.
7 Discussion
In this work, we investigate the influence of parameter initialization scale on the solutions learned by transformers for compositional tasks. We discover a phase diagram of model solutions, where small initialization scales lead to inferential solutions that capture the underlying compositional primitives, while larger initialization scales result in symmetric solutions that memorize mappings. To explain these findings, we analyze the information flow and vector representations within the model, revealing distinct mechanisms for inferential and symmetric solutions. Inferential solutions exhibit low complexity and learn individual mappings for single anchors, while symmetric solutions directly combine anchor information without learning the compositional structure. Building upon these insights, we can predict and experimentally confirm that models with small initialization scales require more training steps to fit data with larger inferential complexity, while models with large initialization scales are less affected by inferential complexity.
Our work opens up promising avenues for future research, such as leveraging Mixture of Experts (MoE) to design networks with different initialization scales for different expert models, and exploring improved training strategies that adapt the model’s focus on enhancing memorization or inference capabilities at different stages of training. It is important to acknowledge the limitations of our current study. The experiments and analyses are based on synthetic data, which allows us to cleanly isolate and characterize the observed phenomena. However, the applicability of our findings to more realistic and complex datasets remains to be validated. In future work, we plan to extend our investigation to real-world datasets and tasks, aiming to bridge the gap between our theoretical understanding and the practical application of these models in solving compositional problems.
Appendix A Experimental Setups
For Fig. 2, Fig. 3, Fig. 4 and Fig. 5, we train the transformer model on a dataset of 900,000 samples, with each input sequence having a length of 9 tokens. The key items in the input sequences are randomly sampled from the range 20-99. The model architecture consists of a varying number of layers, a single attention head, an embedding dimension of 400, a feedforward dimension of 1200, and key and value dimensions of 200 each. The model parameters are initialized using a normal distribution with a mean of 0 and a standard deviation of $(1/d_{\mathrm{in}})^{\gamma}$, where $d_{\mathrm{in}}$ is the input dimension of the parameter and $\gamma$ is the initialization rate. We employ the AdamW optimizer with a learning rate of 1e-5, epsilon of 1e-8, weight decay of 0.01, and $\beta_{1}$ and $\beta_{2}$ values of 0.9 and 0.999, respectively. The model is trained for 210 epochs using a batch size of 2048 and a gradient clipping maximum norm of 1. The learning rate is scheduled using a warm-up period followed by cosine decay, with a warmup period of 10 epochs, a multiplier of 25, a cosine decay number of epochs of 200, and a minimum learning rate of 1e-5. For Fig. 3(a, b), Fig. 4(b), Fig. 5(b, d), we use a two-layer transformer with initialization rate $\gamma=0.5$. For Fig. 3(c), Fig. 4(a, c), Fig. 5(a, c), we use a two-layer transformer with initialization rate $\gamma=0.8$.
For Fig. 6, most of the experimental settings are the same. The difference lies in the dataset size. Specifically, we train the transformer models on a dataset of 100,000 samples. To increase the inferential complexity, we systematically modify the designated target mappings of the following composite anchor pair groups: {(1, 2), (2, 1)}, {(1, 3), (3, 1)}, {(1, 4), (4, 1)}, {(2, 3), (3, 2)}, {(2, 4), (4, 2)}, and {(3, 4), (4, 3)}.
Appendix B Eigenvalue Analysis
Figure 7: Eigenvalues of the covariance matrix of the embedding vectors for different initialization scales. Left: Small initialization scale. Right: Large initialization scale.
To further verify the low complexity bias of word embedding matrices, we visualize the eigenvalues of the covariance matrix of the embedding vectors during training (Fig. 7). For small initialization scales, the covariance matrix has a small number of large eigenvalues, indicating that the model learns a low-dimensional representation of the input tokens that captures their ordinal relationships. This low-dimensional representation facilitates the learning of inferential solutions by aligning with the underlying structure of the single anchor operations.
In contrast, for large initialization scales, the eigenvalues are more evenly distributed, suggesting that the model learns a more distributed representation that does not effectively capture the ordinal structure. This lack of structure in the embedding space hinders the model’s ability to learn the relationships between the input tokens and the underlying single anchor operations, leading to the learning of symmetric solutions.